Will you please answer my question:-7/4a+1/2=-9/10
Answer by malakumar_kos@yahoo.com(315)


Will you please answer my question:-7/4a+1/2=-9/10
-7/4a +1/2 = -9/10
-7/4a = -9/10-1/2
= (-9-5)/10
-7/4a = -14/10
-7.10/-14 = 4a
5 = 4a
a = 5/4